Problem 2. [6 pts]. You bought a stock one year ago for $155 per share and sold i today for $265 per share. It paid a $1.20 per share dividend today. a) what was your realized return? b) How much of the return came from dividend yield and how much came from capital gain? (Show your calculations)
Answer #1

a.Realized return=(End value-Beginning value+Dividends)/Beginning value

=(265-155+1.2)/155

=71.74%(Approx).

b.Dividend yield=Dividend/Beginning value

=(1.2/155)

=0.77%(Approx).

Capital gain=(End price-Beginning value)/Beginning value

=(265-155)/155

=70.97%(Approx).